# symbol/.style= draw=none, every to/.append style= edge node=node [sloped, allow upside down, auto=false]#$1$

### v1

@@ 4.2. Connected Shimura varieties
  • for some faithful representation $\rho:P\to\operatorname{\mathrm{GL}}(V)$ of $P$, the representations $\rho\circ\widetilde{x}$ and $\rho\circ\widetilde{y}$ are conjugate to each other by an element of $\operatorname{\mathrm{GL}}(V)(\mathbb{Q})$ (implicitly in Edixhoven and Yafaev [10], see Remark 3.5 in [24]),
  • for some choice of $\widetilde{y}$, there is an automorphism $\phi:P\to P$ that induces an automorphism of connected Shimura data $(P,X^{+})\to(P,X^{+})$ such that $\widetilde{y}=\phi\circ\widetilde{x}$ (Pink [24]), or
@@
Since there is a Shimura subdatum $(M,X_{M}^{+})$ of $(P,X^{+})$ such that $\widetilde{x}\in X_{M}^{+}$ and since we can always use Lemma 2.2 in [21] to adjust the congruence subgroup, the generalized Hecke orbits of $x$ in the sense of Pink and in the sense of Richard and Yafaev are both contained in $\Sigma_{S_{\mathbb{C}}}(x)$.
@@
Since the definition of a connected mixed Shimura datum/variety in [24] is more restrictive than ours, Conjecture 1.6 in [24] for $S$ and the point $x\in S(\mathbb{C})$ is then a consequence of $\operatorname{\mathrm{WML}}(S_{\mathbb{C}},\Sigma(x),\dim S)$ as defined below in Definition 7.1.
We now want to relate the notion of $\Sigma$-orbits to the definition of structures of finite rank which can be found in Definition 1.5 of [2]. One can easily do this using Lemma 3.7 and the characterization of special subvarieties of modular curves in [9].
@@
A notion of structure of finite rank has also been introduced in [1] in the setting of connected pure Shimura varieties. This definition is different from the one in [2]: in $Y(1)_{\mathbb{C}}^{n}$, a structure of finite rank in the sense of [2] is usually not contained in a structure of finite rank in the sense of [1]. Indeed, it follows from Theorems 2.4 and 2.11 in [30] that a structure of finite rank in $Y(1)_{\mathbb{C}}^{n}$ in the sense of [1] is contained in the union of all special points of $Y(1)_{\mathbb{C}}^{n}$ with a finite union of classical Hecke orbits in $Y(1)_{\mathbb{C}}^{n}$ as defined in [30] and so contained in the union of all special points of $Y(1)_{\mathbb{C}}^{n}$ with a finite union of isogeny classes. On the other hand, as soon as $n\geq 2$ and $\Xi\subset Y(1)(\mathbb{C})$ contains a non-special point $x$, at most finitely many non-special points, and infinitely many pairwise non-isogenous CM points, the structure of finite rank $\overline{\Xi}^{n}$ in $Y(1)_{\mathbb{C}}^{n}$ in the sense of [2] contains points of the form $(x,s,\dots,s)$ ($s\in\Xi$ CM point), which are all non-special and belong to infinitely many isogeny classes.
This shows that, in general, given a $\Sigma$-orbit $\Sigma$ and a connected pure Shimura variety $S$, one cannot find a structure of finite rank in the sense of [1] containing $\Sigma_{S_{\mathbb{C}}}$.
On the other hand, by the above considerations it is clear that, for a connected pure Shimura variety $S$, a structure of finite rank in the sense of [1] in $S_{\mathbb{C}}$ is contained in $\Sigma_{S_{\mathbb{C}}}$ for some $\Sigma$-orbit $\Sigma$.
@@ 5. $\Sigma$-special subvarieties
It follows directly from the definition of being $\Sigma$-special that (3) implies (1). Suppose that (2) holds. Then there exists a distinguished morphism $\chi:Y\to Z\times X$ such that $W\times\{x\}$ is an irreducible component of $\chi(Y)\cap(Z\times\{x\})$. Let $\phi$ and $\psi$ denote the compositions of $\chi$ with the projections to $X$ and $Z$ respectively, then
@@
Hence $W$ is an irreducible component of $\psi(\phi^{-1}(x))$ and (3) holds. Finally, we want to show that (1) implies (2). So suppose that (1) holds. By Lemma 5.3, there exist distinguished morphisms $\phi:Y\to Z^{\prime}$ and $\psi:Y\to Z$ as well as a point $z^{\prime}\in\Sigma_{Z^{\prime}}(x)$ such that $W=\psi(\phi^{-1}(z^{\prime}))$. Unravelling the definition of $\Sigma_{Z^{\prime}}(x)$, we obtain distinguished morphisms $\phi^{\prime}:Y^{\prime}\to X$ and $\psi^{\prime}:Y^{\prime}\to Z^{\prime}$ as well as a point $y^{\prime}\in Y^{\prime}(K)$ such that $\phi^{\prime}$ has finite fibers, $\phi^{\prime}(y^{\prime})=x$, and $\psi^{\prime}(y^{\prime})=z^{\prime}$.
@@
Note that, if $\Sigma=\Sigma(x)$ for some $x\in X(K)$, then the $\Sigma$-special points of $Z$ are exactly the points in $\Sigma_{Z}(x)$, by Lemmas 5.6.
@@
For all results in this section where we assume that (A5) holds, it is actually sufficient to assume that the conclusion of Lemma 5.8 is satisfied, which of course is much weaker than (A5) itself.
@@
We apply Lemma 5.8 to find distinguished morphisms $\psi:Y\to Z$ and $\phi:Y\to X$ as well as a point $x\in X(K)$ such that $W=\psi(\phi^{-1}(x))$ and $\psi$ has finite fibers. Since $W$ is $\Sigma_{Z}$-point-special, there exists $w\in W(K)\cap\Sigma_{Z}$. By Proposition 3.2 and the definition of $\Sigma$-orbit, any $y$ in the non-empty set $(\psi^{-1}(w)\cap\phi^{-1}(x))(K)$ belongs to $\Sigma_{Y}$. It follows that $x\in\Sigma_{X}$ and hence $W$ is $\Sigma$-special. ∎
@@
Part (1) directly follows from Lemma 5.3 and Lemma 6.6 in [3]. Part (2) is clear from Proposition 3.4 and the definitions since the irreducible components of a product are precisely all possible products of an irreducible component of the first factor and an irreducible component of the second factor. Part (3) can be deduced from Proposition 3.2 and Lemma 6.7 in [3] by noting that the proof of Lemma 6.7 in [3] together with Lemma 5.3 implies that irreducible components of pre-images of $\Sigma$-special subvarieties under distinguished morphisms are $\Sigma$-special. Furthermore, every irreducible component of $\phi^{-1}(W)$ of dimension $\dim W$ surjects onto $W$ under $\phi$. ∎
@@
By Lemma 5.8, as $V$ is weakly special, there are distinguished morphisms $\phi:Y\to Z$ and $\psi:Y\to X$ and a point $x\in X(K)$ such that $V=\phi(\psi^{-1}(x))$ and $\phi$ has finite fibers. Because of the latter fact, components of $\phi^{-1}(V)$ have dimension $\leq\dim V$ and clearly we must have $\psi^{-1}(x)\subset\phi^{-1}(V)$. Therefore, also the components of $\psi^{-1}(x)$ have dimension $\leq\dim V$ and there must be one of dimension $\dim V$. This means that $\phi^{-1}(V)$ and $\psi^{-1}(x)$ have a common component of dimension $\dim V$, which we call $V^{\prime}$. Note that $\phi(V^{\prime})=V$ because $\phi(V^{\prime})$ is a weakly special subvariety of the same dimension as $V$. We pick a component $W^{\prime}$ of $\phi|_{V^{\prime}}^{-1}(W)$ such that $\phi(W^{\prime})=W$. This is a component of $\phi^{-1}(W)$ because $\phi$ has finite fibers. Then $W^{\prime}$ is $\Sigma$-special and the same holds for $\{x\}$, by Proposition 5.11. Another application of the same proposition gives the desired claim. ∎
@@ 6.1. Semiabelian varieties
Let $G$ be a semiabelian variety. We want to see how the concepts of a $\Sigma$-special subvariety and of a $\Sigma_{G}$-point-special subvariety of $G$ are related in this context. Since the distinguished category of semiabelian varieties satisfies (A5), we know already that a $\Sigma_{G}$-point-special subvariety of $G$ is $\Sigma$-special.
@@
We can and will assume that $x$ does not belong to any proper special subvariety of $G_{0}$ as done in Subsection 4.1. Let $W$ be a $\Sigma$-special subvariety of $G$. Then, $W$ is a component of $\psi(\phi^{-1}(x))$ for some distinguished morphisms $\psi:Y\to G$ and $\phi:Y\to G_{0}$. Since $x$ does not belong to any proper special subvariety of $G_{0}$, we have that $\phi$ is surjective. We may assume as in Subsection 4.1 that $\phi$ is a homomorphism of algebraic groups. Since $Y$ is almost-split, there exist an abelian variety $B$ over $K$, an integer $n\geq 0$, and an isogeny $\omega:B\times\mathbb{G}^{n}_{m,K}\to Y$. The kernel of $\phi\circ\omega$ is of the form $H\times H^{\prime}$ for algebraic subgroups $H$ of $B$ and $H^{\prime}$ of $\mathbb{G}^{n}_{m,K}$ respectively. It then follows from Poincaré’s complete reducibility theorem (see, e.g., [20, Theorem 1, p. 173]) that there is a semiabelian subvariety of $B\times\mathbb{G}^{n}_{m,K}$ such that $\phi\circ\omega$ restricts to an isogeny from this semiabelian subvariety to $G_{0}$. Therefore, there also exists a semiabelian subvariety of $Y$ such that $\phi$ restricts to an isogeny from this semiabelian subvariety to $G_{0}$. By Lemma 7.3/5 on p. 180 of [4], we can conclude that there exists a homomorphism of algebraic groups $\phi^{\prime}:G_{0}\to Y$ such that $\chi=\phi\circ\phi^{\prime}$ is the multiplication by some positive integer $d$. Then, $\chi^{-1}(x)$ is a finite set of points which lie in $\Sigma_{G_{0}}$ by Proposition 5.11 and some component of $\phi^{-1}(x)$ must contain a point of $\phi^{\prime}(\chi^{-1}(x))$ which must be a point of $\Sigma_{Y}$ by the same Proposition 5.11. But the components of $\phi^{-1}(x)$ are all translates of each other by torsion points of $Y$ and since translating by a torsion point is a distinguished morphism, each of them contains a point of $\Sigma_{Y}$ by Proposition 5.11. Applying Proposition 5.11 once more, we deduce that any component of $\psi(\phi^{-1}(x))$ contains a point of $\Sigma_{G}$. ∎
@@
We now see that, if the almost-split condition is dropped, meaning we consider the distinguished category of semiabelian varieties, the statement of the above proposition cannot hold. Let $G_{0}$ be an arbitrary semiabelian variety and let $x_{0}\in G_{0}(K)$. Let $E$ be an elliptic curve with $\mathrm{End}(E)\simeq\mathbb{Z}$ and let $y_{0}\in E(K)$ be a non-torsion point. For any abelian variety $A$, we denote its dual abelian variety by $\widehat{A}$. The extensions of $A$ by $\mathbb{G}_{m,K}^{n}$ over $K$ are classified by $\widehat{A}^{n}(K)$, see Section 1.1 of [16]. Let $G$ be an extension of $E$ by $\mathbb{G}_{m,K}$ that corresponds to $g\in\widehat{E}(K)$. We have the canonical projection $\pi:G\to E$ and $\pi^{-1}(y_{0})$ is a $\Sigma(x_{0},y_{0})$-special subvariety of $G$. We are going to see that, for “most choices” of $g$, $\pi^{-1}(y_{0})$ is not $\Sigma_{G}(x_{0},y_{0})$-point-special. Suppose that $\phi:G_{0}^{\prime}\to G_{0}\times E$ is a homomorphism of semiabelian varieties with finite kernel. Let $A_{0}$ and $A_{0}^{\prime}$ denote the abelian quotients of $G_{0}$ and $G_{0}^{\prime}$ respectively and let $g_{0}\in\widehat{A_{0}}^{m}(K)$ and $g_{0}^{\prime}\in\widehat{A_{0}^{\prime}}^{m^{\prime}}(K)$ denote the elements corresponding to $G_{0}$ and $G_{0}^{\prime}$ respectively. It follows from Section 1 of [16] that $g_{0}^{\prime}\in(\Gamma^{\prime\mathrm{div}})^{m^{\prime}}$ where $\Gamma^{\prime}=\left\langle\mathrm{Hom}(\widehat{A_{0}\times E},\widehat{A_{0}^{\prime}})\cdot(g_{0},0_{\widehat{E}})\right\rangle$ is a finitely generated subgroup of $\widehat{A_{0}^{\prime}}(K)$. Recall that $\Gamma^{\prime\mathrm{div}}$ indicates the division group of $\Gamma^{\prime}$ in $\widehat{A_{0}^{\prime}}(K)$. Suppose now that $\psi:G_{0}^{\prime}\to G$ is an arbitrary homomorphism of semiabelian varieties. It follows again from Section 1 of [16] that either $\pi\circ\psi$ is trivial or $\chi(g)\in\Gamma^{\prime\mathrm{div}}$ for some homomorphism $\chi:\widehat{E}\to\widehat{A_{0}^{\prime}}$ with finite kernel. Since $\mathrm{End}(\widehat{E})\simeq\mathrm{End}(E)\simeq\mathbb{Z}$, we find in the latter case that $g\in\Gamma^{\mathrm{div}}$ where $\Gamma=\left\langle\mathrm{Hom}(\widehat{A_{0}\times E},\widehat{E})\cdot(g_{0},0_{\widehat{E}})\right\rangle$. Hence, if we choose $G$ such that $g\not\in\Gamma^{\mathrm{div}}$ (this can be done by work of Frey and Jarden [11]), then every $\Sigma(x_{0},y_{0})$-special point of $G$ must project to a torsion point of $E$ under $\pi$. It follows that $\pi^{-1}(y_{0})$ cannot be $\Sigma_{G}(x_{0},y_{0})$-point-special.
@@
In the distinguished category of semiabelian varieties, the following holds: let $G_{0}$ be a semiabelian variety, let $\Sigma$ be the $\Sigma$-orbit of $x\in G_{0}(K)$, and let $G$ be another fixed semiabelian variety. There exists a $\Sigma$-orbit Σsuperscriptnormal-$\Sigma^{\prime}$ such that every $\Sigma$-special subvariety of $G$ is Σsuperscriptnormal-$\Sigma^{\prime}$-special and every Σsuperscriptnormal-$\Sigma^{\prime}$-special subvariety of $G$ is $\Sigma^{\prime}_{G}$-point-special.

### v2

@@ 4.2. Connected pure Shimura varieties
  • for some faithful representation $\rho:P\to\operatorname{\mathrm{GL}}(V)$ of $P$, the representations $\rho\circ\widetilde{x}$ and $\rho\circ\widetilde{y}$ are conjugate to each other by an element of $\operatorname{\mathrm{GL}}(V)(\mathbb{Q})$ (implicitly in Edixhoven and Yafaev [12], see Remark 3.5 in [27] and use the correspondence between Hodge structures and representations of the Deligne torus),
  • for some choice of $\widetilde{y}$, there is an automorphism $\phi:P\to P$ such that $\widetilde{y}=\phi\circ\widetilde{x}$ (Pink [27]), or
@@
Since there is a Shimura subdatum $(M,X_{M}^{+})$ of $(P,X^{+})$ such that $\widetilde{x}_{0}\in X_{M}^{+}$ and since we can always use Lemma 2.2 in [24] to adjust the congruence subgroup, the generalized Hecke orbits of $x$ in the sense of Pink and in the sense of Richard and Yafaev are both contained in $\Sigma_{S_{\mathbb{C}}}(x)$.
@@
We now want to relate the notion of $\Sigma$-orbits to the definition of structures of finite rank which can be found in Definition 1.5 of [2]. One can easily do this using Lemma 3.7 and the characterization of special subvarieties of products of modular curves in [11].
@@
A notion of structure of finite rank has also been introduced in [1] in the setting of connected pure Shimura varieties. This definition is different from the one in [2]: in $Y(1)_{\mathbb{C}}^{n}$, a structure of finite rank in the sense of [2] is usually not contained in a structure of finite rank in the sense of [1]. Indeed, it can be deduced from Theorems 2.4 and 2.11 in [32] that a structure of finite rank in $Y(1)_{\mathbb{C}}^{n}$ in the sense of [1] is contained in the union of all special points of $Y(1)_{\mathbb{C}}^{n}$ with a finite union of classical Hecke orbits in $Y(1)_{\mathbb{C}}^{n}$ as defined in [32] and so contained in the union of all special points of $Y(1)_{\mathbb{C}}^{n}$ with a finite union of isogeny classes. On the other hand, as soon as $n\geq 2$ and $\Xi\subset Y(1)(\mathbb{C})$ contains a non-special point $x$, at most finitely many non-special points, and infinitely many CM points that belong to pairwise distinct isogeny classes, the structure of finite rank $\overline{\Xi}^{n}$ in $Y(1)_{\mathbb{C}}^{n}$ in the sense of [2] contains points of the form $(x,s,\dots,s)$ ($s\in\Xi$ CM point), which are all non-special and belong to infinitely many isogeny classes.
This shows that, in general, given a $\Sigma$-orbit $\Sigma(x)$ and a connected pure Shimura variety $S$, one cannot find a structure of finite rank in the sense of [1] containing $\Sigma_{S_{\mathbb{C}}}(x)$.
On the other hand, by the above considerations it is clear that, for a connected pure Shimura variety $S$, a structure of finite rank in the sense of [1] in $S_{\mathbb{C}}$ is contained in $\Sigma_{S_{\mathbb{C}}}(x)$ for some $\Sigma$-orbit $\Sigma(x)$.
Finally, let us see that our notion of $\Sigma$-orbit is equivalent to the notion of hybrid orbit by Richard and Yafaev introduced in Definition 6 in [31]. Suppose we have a Hodge generic point $x\in X^{+}_{M}$ for a connected pure Shimura variety $\mathcal{F}_{\mathrm{pSv}}(M,X^{+}_{M},\Gamma_{M})$. The hybrid Hecke orbit of $x$ in $\mathcal{F}_{\mathrm{pSv}}(P,X^{+},\Gamma)$ is the image of the set of all $x^{\prime}\in X^{+}$ such that $\phi(x^{\text{ad}})=x^{\prime\text{ad}}$ for a homomorphism $\phi:M^{\text{ad}}\to M_{x^{\prime}}^{\text{ad}}$ (that therefore induces a Shimura morphism $(M^{\text{ad}},X_{M}^{+,\text{ad}})\to(M_{x^{\prime}}^{\text{ad}},X_{x^{\prime}}^{+,\text{ad}})$), where $M_{x^{\prime}}\subset P$ denotes the Mumford-Tate group of $x^{\prime}$ and $X^{+}_{x^{\prime}}=M_{x^{\prime}}(\mathbb{R})^{+}\cdot x^{\prime}$. The image in $\Gamma\backslash X^{+}$ of any such $x^{\prime}$ will be in the $\Sigma$-orbit of the image of $x$ under $X^{+}_{M}\to\Gamma_{M}\backslash X^{+}_{M}$ according to our definition because of Propositions 3.2 and 3.3 since:
  • $(M,X^{+}_{M})\to(M^{\text{ad}},X_{M}^{+,\text{ad}})$ is a Shimura immersion and
  • $(M_{x^{\prime}},X^{+}_{x^{\prime}})\to(M_{x^{\prime}}^{\text{ad}},X_{x^{\prime}}^{+,\text{ad}})$ is a Shimura immersion.
Vice versa, suppose that $x^{\prime}$ is a pre-image in $X^{+}$ of a point of our $\Sigma$-orbit of $x$ in $\Gamma\backslash X^{+}$. It can be deduced from Theorem 7.1 in [31] and from Lemma 3.7 that $x^{\prime}$ also belongs to the hybrid Hecke orbit of $x$.
@@ 5. $\Sigma$-special subvarieties
It follows directly from the definition of being $\Sigma$-special that (3) implies (1).
Suppose that (2) holds. Then there exists a distinguished morphism $\chi:Y\to Z\times X$ such that $W\times\{x\}$ is an irreducible component of $\chi(Y)\cap(Z\times\{x\})$. Let $\phi$ and $\psi$ denote the compositions of $\chi$ with the projections to $X$ and $Z$ respectively, then
@@
Hence $W$ is an irreducible component of $\psi(\phi^{-1}(x))$ and (3) holds.
Finally, we want to show that (1) implies (2). So suppose that (1) holds. By Lemma 5.3, there exist distinguished morphisms $\phi:Y\to Z^{\prime}$ and $\psi:Y\to Z$ as well as a point $z^{\prime}\in\Sigma_{Z^{\prime}}(x)$ such that $W=\psi(\phi^{-1}(z^{\prime}))$. Unravelling the definition of $\Sigma_{Z^{\prime}}(x)$, we obtain distinguished morphisms $\phi^{\prime}:Y^{\prime}\to X$ and $\psi^{\prime}:Y^{\prime}\to Z^{\prime}$ as well as a point $y^{\prime}\in Y^{\prime}(K)$ such that $\phi^{\prime}$ has finite fibers, $\phi^{\prime}(y^{\prime})=x$, and $\psi^{\prime}(y^{\prime})=z^{\prime}$.
@@
Note that, if $\Sigma=\Sigma(x)$ for some $x\in X(K)$, then the $\Sigma$-special points of $Z$ are exactly the points in $\Sigma_{Z}(x)$, by Lemma 5.6.
@@
For all results in this section where we assume that (A5) holds, it is actually sufficient to assume that the conclusion of Lemma 5.8 is satisfied, which of course is much weaker than (A5) itself. For instance, in the distinguished category $\mathfrak{C}_{\mathrm{add}}$ whose objects are the powers of the additive group $\mathbb{G}_{a,K}$, (A5) does not hold (see Section 9 of [4]), but the conclusion of Lemma 5.8 does.
@@
We apply Lemma 5.8 to find distinguished morphisms $\psi:Y\to Z$ and $\phi:Y\to X$ as well as a point $x\in X(K)$ such that $W=\psi(\phi^{-1}(x))$ and $\psi$ has finite fibers. Since $W$ is $\Sigma_{Z}$-point-special, there exists $w\in W(K)\cap\Sigma_{Z}$. By Proposition 3.2 and the definition of a $\Sigma$-orbit, any $y$ in the non-empty set $(\psi^{-1}(w)\cap\phi^{-1}(x))(K)$ belongs to $\Sigma_{Y}$. It follows that $x\in\Sigma_{X}$ and hence $W$ is $\Sigma$-special. ∎
@@
Part (1) directly follows from Lemma 5.3 and Lemma 6.7 in [4].
Part (2) is clear from Proposition 3.4 and the definitions since the irreducible components of a product are precisely all possible products of an irreducible component of the first factor and an irreducible component of the second factor.
Part (3) can be deduced from Proposition 3.2 and Lemma 6.7 in [4] by noting that the proof of Lemma 6.7 in [4] together with Lemma 5.3 implies that irreducible components of pre-images of $\Sigma$-special subvarieties under distinguished morphisms are $\Sigma$-special. Furthermore, every irreducible component of $\phi^{-1}(W)$ of dimension $\dim W$ surjects onto $W$ under $\phi$. ∎
@@
By Lemma 5.8, as $V$ is weakly special, there are distinguished morphisms $\phi:Y\to Z$ and $\psi:Y\to X$ and a point $x\in X(K)$ such that $V=\phi(\psi^{-1}(x))$ and $\phi$ has finite fibers. Because of the latter fact, components of $\phi^{-1}(V)$ have dimension $\leq\dim V$ and clearly we must have $\psi^{-1}(x)\subset\phi^{-1}(V)$. Therefore, also the components of $\psi^{-1}(x)$ have dimension $\leq\dim V$ and there must be one of dimension $\dim V$. This means that $\phi^{-1}(V)$ and $\psi^{-1}(x)$ have a common component of dimension $\dim V$, which we call $V^{\prime}$. Note that $\phi(V^{\prime})=V$ because $\phi(V^{\prime})$ is a weakly special subvariety of the same dimension as $V$.
We pick a component $W^{\prime}$ of $\phi|_{V^{\prime}}^{-1}(W)$ such that $\phi(W^{\prime})=W$. This is a component of $\phi^{-1}(W)$ because $\phi$ has finite fibers. Then $W^{\prime}$ is $\Sigma$-special and the same holds for $\{x\}$, by Proposition 5.11. Another application of the same proposition gives the desired claim. ∎
@@ 6.1. Semiabelian varieties
Let $G$ be a semiabelian variety. We want to see how the concepts of a $\Sigma$-special subvariety and of a $\Sigma_{G}$-point-special subvariety of $G$ are related in this context. Since the distinguished category of semiabelian varieties satisfies (A5) by Theorem 9.3 in [4], we know already that a $\Sigma_{G}$-point-special subvariety of $G$ is $\Sigma$-special.
@@
We can and will assume that $x$ does not belong to any proper special subvariety of $G_{0}$ as done in Subsection 4.1.
Let $W$ be a $\Sigma$-special subvariety of $G$. Then, $W$ is a component of $\psi(\phi^{-1}(x))$ for some distinguished morphisms $\psi:Y\to G$ and $\phi:Y\to G_{0}$. Since $x$ does not belong to any proper special subvariety of $G_{0}$, we have that $\phi$ is surjective. We may assume as in Subsection 4.1 that $\phi$ is a homomorphism of algebraic groups. Since $Y$ is almost-split, there exist an abelian variety $B$ over $K$, an integer $n\geq 0$, and an isogeny $\omega:B\times\mathbb{G}^{n}_{m,K}\to Y$. The kernel of $\phi\circ\omega$ is of the form $H\times H^{\prime}$ for algebraic subgroups $H$ of $B$ and $H^{\prime}$ of $\mathbb{G}^{n}_{m,K}$ respectively. It then follows from Poincaré’s complete reducibility theorem (see, e.g., [23, Theorem 1, p. 173]) that there is a semiabelian subvariety of $B\times\mathbb{G}^{n}_{m,K}$ such that $\phi\circ\omega$ restricts to an isogeny from this semiabelian subvariety to $G_{0}$. Therefore, there also exists a semiabelian subvariety of $Y$ such that $\phi$ restricts to an isogeny from this semiabelian subvariety to $G_{0}$. By Lemma 7.3/5 on p. 180 of [6], we can conclude that there exists a homomorphism of algebraic groups $\phi^{\prime}:G_{0}\to Y$ such that $\chi=\phi\circ\phi^{\prime}$ is the multiplication by some positive integer $d$. Then, $\chi^{-1}(x)$ is a finite set of points which lie in $\Sigma_{G_{0}}$ by Proposition 5.11 and some component of $\phi^{-1}(x)$ must contain a point of $\phi^{\prime}(\chi^{-1}(x))$ which must be a point of $\Sigma_{Y}$ by the same Proposition 5.11. But the components of $\phi^{-1}(x)$ are all translates of each other by torsion points of $Y$ and since translating by a torsion point is a distinguished morphism, each of them contains a point of $\Sigma_{Y}$ by Proposition 5.11. Applying Proposition 5.11 once more, we deduce that any component of $\psi(\phi^{-1}(x))$ contains a point of $\Sigma_{G}$. ∎
@@
We now see that, if the almost-split condition is dropped, meaning we consider the distinguished category of semiabelian varieties, the statement of the above proposition cannot hold.
Let $G_{0}$ be an arbitrary semiabelian variety and let $x_{0}\in G_{0}(K)$. Let $E$ be an elliptic curve with $\mathrm{End}(E)\simeq\mathbb{Z}$ and let $y_{0}\in E(K)$ be a non-torsion point.
For any abelian variety $A$, we denote its dual abelian variety by $\widehat{A}$. The extensions of $A$ by $\mathbb{G}_{m,K}^{n}$ over $K$ are classified by $\widehat{A}^{n}(K)$, see Section 1.1 of [19]. Let $G$ be an extension of $E$ by $\mathbb{G}_{m,K}$ that corresponds to $g\in\widehat{E}(K)$.
We have the canonical projection $\pi:G\to E$ and $\pi^{-1}(y_{0})$ is a $\Sigma(x_{0},y_{0})$-special subvariety of $G$. We are going to see that, for “most choices” of $g$, $\pi^{-1}(y_{0})$ is not $\Sigma_{G}(x_{0},y_{0})$-point-special.
Suppose that $\phi:G_{0}^{\prime}\to G_{0}\times E$ is a homomorphism of semiabelian varieties with finite kernel. Let $A_{0}$ and $A_{0}^{\prime}$ denote the abelian quotients of $G_{0}$ and $G_{0}^{\prime}$ respectively and let $g_{0}\in\widehat{A_{0}}^{m}(K)$ and $g_{0}^{\prime}\in\widehat{A_{0}^{\prime}}^{m^{\prime}}(K)$ denote the elements corresponding to $G_{0}$ and $G_{0}^{\prime}$ respectively. It follows from Section 1 of [19] that $g_{0}^{\prime}\in(\Gamma^{\prime\mathrm{div}})^{m^{\prime}}$ where $\Gamma^{\prime}=\left\langle\mathrm{Hom}(\widehat{A_{0}\times E},\widehat{A_{0}^{\prime}})\cdot(g_{0},0_{\widehat{E}})\right\rangle$ is a finitely generated subgroup of $\widehat{A_{0}^{\prime}}(K)$. Recall that $\Gamma^{\prime\mathrm{div}}$ indicates the division group of $\Gamma^{\prime}$ in $\widehat{A_{0}^{\prime}}(K)$. Suppose now that $\psi:G_{0}^{\prime}\to G$ is an arbitrary homomorphism of semiabelian varieties. It follows again from Section 1 of [19] that either $\pi\circ\psi$ is trivial or $\chi(g)\in\Gamma^{\prime\mathrm{div}}$ for some homomorphism $\chi:\widehat{E}\to\widehat{A_{0}^{\prime}}$ with finite kernel. Since $\mathrm{End}(\widehat{E})\simeq\mathrm{End}(E)\simeq\mathbb{Z}$, we find in the latter case that $g\in\Gamma^{\mathrm{div}}$ where $\Gamma=\left\langle\mathrm{Hom}(\widehat{A_{0}\times E},\widehat{E})\cdot(g_{0},0_{\widehat{E}})\right\rangle$.
Hence, if we choose $G$ such that $g\not\in\Gamma^{\mathrm{div}}$ (this can be done by work of Frey and Jarden [13]), then every $\Sigma(x_{0},y_{0})$-special point of $G$ must project to a torsion point of $E$ under $\pi$. It follows that $\pi^{-1}(y_{0})$ cannot be $\Sigma_{G}(x_{0},y_{0})$-point-special.
@@
In the distinguished category of semiabelian varieties, the following holds: let $G_{0}$ be a semiabelian variety, let $\Sigma$ be the $\Sigma$-orbit of $x\in G_{0}(K)$, and let $G$ be another fixed semiabelian variety. There exist $\gamma_{1}^{\prime},\ldots,\gamma_{n}^{\prime}\in G(K)$ such that, for $\Sigma^{\prime}=\Sigma(x,\gamma_{1}^{\prime},\ldots,\gamma_{n}^{\prime})$, we have that every $\Sigma^{\prime}$-special subvariety of $G$ is $\Sigma^{\prime}_{G}$-point-special.